# Do LLMs write like humans? Variation in grammatical and rhetorical styles

### v1

@@ 3.3 Distinguishing individual LLMs
3.4 Generalization across corpora
When each pairwise random forest was used to classify arXiv preprints from the M4 corpus (Wang et al., 2024), accuracy dropped significantly. Models trained on instruction-tuned LLMs were able to classify M4’s GPT-3.5 output with greater-than-chance accuracy, but models trained on the Llama base outputs attained only 50% accuracy, equal to random guessing. Full results are presented in Tables LABEL:tbl:hape-v-arxiv and LABEL:tbl:coca-v-arxiv. These results demonstrate that instruction-tuned LLMs do have features in common that permit their classification, but that generalizability across LLMs or to different registers of text is difficult.
4 Discussion
This study identifies salient differences both between human and LLM-generated texts and among various models. The features that distinguish between humans and different LLMs include present participial clauses, ‘that’ clauses as sentence subject, passive voice, and nominalizations, to name a few. These findings corroborate other research that points out the ways these features produce informationally dense prose (Markey et al., 2024). Other research links these features to increased lexical diversity in generated text, as well as human judgments of linguistic mastery (Herbold et al., 2023). Lastly, Jiang and Hyland (2024) found that ChatGPT-4 text evidences more nominalizations, and fewer human subjects and epistemic stance markers, findings we see reproduced in our list of distinguishing features.
A second major finding of this research is the apparently central role of instruction tuning in creating these discrepancies between human and model general texts. While we do not have access to untuned versions of GPT, comparisons between Llama’s base and tuned models emphasize the degree to which instruction tuning pushes models to produce text that reads unlike a human. This suggests that differences in style are not simply due to the selection of texts for training the base models, but due to human preferences in instruction tuning, either deliberately or as an unintended consequence of other goals. To our knowledge, this is not discussed elsewhere in similar research.
@@ 4 Discussion
As LLMs are increasingly put to work completing diverse writing tasks, these results suggest a notable misalignment between generated texts and the contexts in which we put them to use. This is another way of saying that LLMs do not vary their linguistic output in response to contextual factors in ways similar to humans. For those proficient in a genre—think, a therapist collating notes on a patient or, say, a college graduate writing thank-you notes to friends and family—this misalignment may not be a problem. Previous experience, in this case, guides a current sense of what a particular document should look like in order to be successful. For experts, then, LLMs appear a worthwhile productivity tool, suitable so long as they lend their expertise to further shaping the output.
For learners, though, LLMs appear problematic. Of course, using LLMs to learn more about a concept, or help generate ideas, is one thing. Using output in a text is a different matter, one that may affect a student’s learning trajectories. In this case, students offload the important cognitive labor of shaping a text for a particular audience and purpose to the LLM. Never mind that LLMs do not appear to write like humans—when students offload writing work, they offload opportunities to learn how to write IMRaD articles, client reports, executive summaries, investment pitches, etc. When LLMs are used in the classroom as writing tools, instructors of all levels and disciplines need to help students see both the shortcomings of the generated text and avenues for improvement. LLMs are not bad, either technically or morally—it is only that instructors must help inculcate in students the critical perspective of the expert to know what’s working and what isn’t.
@@ Acknowledgments
Data availability
The Human-AI Parallel English (HAP-E) Corpus described in Section 2.1 is available at:
The Corpus of Contemporary American English (COCA) can be obtained at https://www.corpusdata.org/. Due to the licensing agreement, we cannot redistribute the text, but Biber features for the subset analyzed above can be found at https://huggingface.co/datasets/browndw/coca-ai-parallel-corpus-biber
Appendix A Appendix
Table 2: Lexical, grammatical, and rhetorical features detected by pseudobibeR, adapted from Biber (1988, Appendix II).
Feature Description Examples
Past tense Verbs in past tense I ran far.
Perfect aspect Verbs in perfect aspect I have written this sentence.
Present tense Verbs in present tense. the duck walks quickly
Place adverbials Adverbs and adverbial phrases describing place above, beside, outdoors
First-person pronouns I, we, our
Second-person pronouns you, your
Third-person pronouns (excluding it) he, she, they, their
Pronoun it it, its, itself
Demonstrative pronouns Pronouns replacing nouns That is an example sentence.
Indefinite pronouns anybody, nothing, someone
Pro-verb ‘do’
‘Wh-’ questions Direct who, what, when, where, and why questions When are you leaving?
Nominalizations Nouns ending in -tion, -ment, etc. justification, development, robustness
Gerunds Participial forms functioning as nouns
Other nouns All other nouns
Agentless passives Agentless passive voice The model was fitted
‘By-’ passives Passive voice with agent The task was done by Steve
‘Be’ as main verb Use of “be” forms as main verb
Existential ‘there’ “There” used to assert something exists There is a feature in this sentence
‘That’ verb complements I said that he went
‘That’ adjective complements I’m glad that you like it
‘Wh-’ clauses Clauses beginning with ‘wh-’ words (who, what, when, …) I believed what he told me
Infinitives Uninflected verb preceded by to she tried to explain
Present participial clauses Adverbial clauses used as present participles Stuffing his mouth with cookies, Joe ran out the door
Past participial clauses Adverbial clauses used as past participles Built in a single week, the house would stand for fifty years
Past participial postnominal Reduced relative past participial clauses The solution produced by this process
Present participial postnominal Reduced relative present participial clauses The event causing this decline
‘That’ clauses as subject ‘That’ relative clauses in subject position the dog that bit me
‘That’ clauses as object ‘That’ relative clauses in object position the dog that I saw
‘Wh-’ relatives as subject ‘Wh-’ relatives in subject position the man who likes popcorn
‘Wh-’ relatives as object ‘Wh-’ relatives in object position the man who Sally likes
Pied-piping relative clauses Relative clauses moved in sentence by ‘wh-’ questions the manner in which he was told
Sentence relatives Bob likes fried mangoes, which is the most disgusting thing I’ve ever heard of
Because Causative adverbial subordinator because
Though Concessive adverbial subordinators although, though
If, unless Conditional adverbial subordinators if, unless
Other adverbial subordinators since, while, whereas
Prepositional phrases
Attributive adjectives the big horse
Predicative adjectives the horse is big
Adverbs Total adverbs
Type-token ratio Type-token ratio, including punctuation
Mean word length Average word length in characters, excluding punctuation
Conjuncts consequently, furthermore, however
Downtoners barely, nearly, slightly
Hedges at about, something like, almost
Amplifiers absolutely, extremely, perfectly
Emphatics a lot, for sure, really
Discourse particles sentence-initial well, now, anyway
Demonstratives “That”, “this”, “these”, or “those” used as determiners That is the feature
Possibility modals can, may, might, could
Necessity modals ought, should, must
Predictive modals will, would, shall
Public verbs Verbs indicating speaking or announcing assert, declare, mention, predict, swear
Private verbs Verbs expressing a personal intellectual state assume, believe, doubt, presume, understand
Suasive verbs Verbs indicating persuasion command, insist, propose
‘Seem’ and ‘appear’ Used as verbs
Contractions can’t, won’t
‘That’ deletion ‘That’ omitted as subordinator I think [that] he went
Stranded prepositions Preposition at the end of sentence or clause rather than before its object she is the candidate that I was thinking of
Split infinitives Infinitive with adverb between to and the verb he wants to convincingly prove that…
Split auxiliaries they were apparently shown to…
Phrasal co-ordination Pairs of nouns, verbs, adjectives, or adverbs connected by a coordinating conjunction The nouns and verbs are coordinated
Clausal co-ordination Sentence clauses coordinated by a conjunction The sentence was long, but I read it anyway
Synthetic negation no answer is good enough
Analytic negation that isn’t good enough
Table 3: Biber features distinguishing human- and LLM-written text in HAP-E Rate per 1,000 tokens; LLM rates relative to human Chunk 2. Blue indicates higher-than-human usage, red indicates lower. Ordered by feature importance for the seven-class classification task.
Human GPT Llama 3 Instruct Llama 3 Base
Feature Chunk 1 Chunk 2 GPT-4o Mini GPT-4o 8B 70B 8B 70B Importance
Present participial clauses 1.7 1.7 481% 527% 224% 261% 94% 102% 1,715.1
‘That’ clauses as subject 2.1 2.1 331% 263% 180% 173% 64% 68% 1,387.2
Mean word length 4.5 4.4 114% 116% 101% 103% 99% 100% 1,070.4
Adverbs 67.9 71.8 86% 82% 73% 75% 102% 102% 916.0
Nominalizations 14.6 14.6 209% 214% 145% 151% 88% 91% 808.9
Phrasal co-ordination 6.7 6.1 144% 194% 187% 170% 92% 97% 776.0
Clausal co-ordination 11.2 12.4 63% 59% 141% 127% 120% 116% 727.3
Other nouns 254.9 240.6 97% 103% 91% 95% 91% 94% 725.5
Downtoners 1.9 1.9 155% 118% 60% 57% 68% 73% 712.8
Attributive adjectives 46.7 43.8 140% 150% 100% 104% 79% 83% 677.7
Demonstratives 6.2 6.5 137% 133% 77% 80% 75% 80% 675.9
Agentless passives 7.7 7.8 51% 53% 96% 89% 101% 98% 664.4
Past participial postnominal 1.6 1.5 257% 235% 75% 75% 129% 131% 633.7
Infinitives 15.9 16.5 87% 83% 140% 132% 120% 113% 614.9
Prepositional phrases 101.0 98.0 118% 118% 100% 100% 87% 90% 581.2
Emphatics 8.6 9.2 75% 68% 97% 98% 76% 75% 545.9
Existential ‘there’ 1.9 2.1 59% 71% 42% 42% 108% 109% 527.3
Possibility modals 5.3 5.7 144% 104% 111% 116% 99% 99% 521.1
Past tense 37.5 41.9 77% 83% 91% 83% 115% 111% 519.0
‘That’ verb complements 2.1 2.5 70% 55% 144% 114% 159% 147% 503.4
Contractions 16.9 18.1 63% 60% 141% 139% 142% 129% 497.7
Other adverbial subordinators 5.9 6.2 136% 118% 114% 107% 82% 89% 496.6
‘Be’ as main verb 29.0 30.0 61% 63% 107% 100% 108% 101% 494.3
Public verbs 5.8 6.8 53% 63% 65% 67% 119% 112% 466.1
Present participial postnominal 1.4 1.3 293% 243% 125% 124% 101% 113% 451.0
Predicative adjectives 5.8 6.2 84% 90% 140% 131% 169% 154% 450.8
Suasive verbs 2.9 3.2 98% 116% 89% 90% 141% 137% 449.6
Analytic negation 8.5 9.7 73% 61% 80% 78% 113% 107% 439.6
Third-person pronouns 26.2 29.8 91% 91% 108% 104% 123% 120% 436.8
Private verbs 16.0 18.1 91% 85% 122% 113% 128% 126% 435.9
Perfect aspect 7.5 7.2 62% 60% 121% 111% 96% 92% 434.2
Pro-verb ‘do’ 2.9 3.2 25% 26% 60% 59% 115% 109% 430.8
Present tense 57.2 58.3 82% 76% 99% 100% 119% 116% 418.9
Amplifiers 2.0 2.1 63% 46% 40% 45% 85% 85% 412.6
Pronoun ‘it’ 12.5 13.2 88% 92% 105% 105% 108% 105% 412.0
Gerunds 3.0 3.0 156% 152% 119% 124% 83% 86% 402.9
First-person pronouns 34.0 35.3 81% 62% 111% 108% 136% 127% 393.4
Because 1.2 1.5 19% 20% 38% 38% 121% 105% 387.9
Split auxiliaries 3.4 3.4 91% 77% 126% 118% 98% 92% 379.4
Time adverbials 3.7 3.7 77% 79% 68% 70% 115% 119% 366.0
Though, although 0.5 0.5 65% 129% 21% 28% 83% 90% 361.2
Place adverbials 3.4 3.4 146% 144% 101% 99% 118% 128% 353.9
Demonstrative pronouns 5.5 6.1 55% 50% 71% 76% 99% 97% 350.8
Predictive modals 5.1 5.6 72% 57% 95% 106% 118% 111% 345.4
Indefinite pronouns 3.2 3.3 73% 77% 84% 83% 156% 149% 323.5
by-passives 0.9 0.9 56% 63% 118% 107% 113% 103% 314.9
Synthetic negation 1.2 1.3 36% 51% 36% 36% 93% 91% 311.8
‘Wh-’ questions 1.3 1.4 89% 56% 141% 137% 140% 132% 306.1
If, unless 2.2 2.6 78% 60% 58% 62% 117% 111% 305.1
‘Seem’ and ‘appear’ 0.7 0.7 131% 179% 140% 128% 99% 105% 290.3
Necessity modals 1.1 1.3 105% 78% 51% 54% 143% 137% 286.7
Conjuncts 2.1 2.3 110% 96% 97% 116% 99% 104% 281.2
Second-person pronouns 14.9 15.7 63% 52% 77% 81% 118% 110% 277.3
Stranded prepositions 0.8 0.9 66% 66% 111% 114% 87% 81% 266.5
‘Wh-’ relatives as subject 2.3 2.1 70% 66% 109% 102% 106% 103% 259.1
Sentence relatives 1.0 1.0 50% 51% 104% 117% 101% 92% 251.7
‘Wh-’ clauses 1.7 1.9 75% 66% 90% 78% 126% 122% 247.1
Discourse particles 1.1 1.0 60% 60% 89% 91% 146% 140% 181.9
Hedges 1.1 1.3 50% 63% 62% 67% 89% 92% 181.1
Past participial clauses 0.3 0.3 273% 307% 158% 150% 59% 71% 171.7
‘That’ deletion 0.7 0.8 75% 66% 75% 73% 134% 151% 169.8
Pied-piping relative clauses 0.6 0.6 59% 56% 38% 43% 50% 54% 164.5
‘That’ clauses as object 0.6 0.6 65% 56% 66% 60% 93% 90% 118.3
‘That’ adjective complements 0.3 0.4 80% 51% 121% 106% 148% 139% 107.6
Split infinitives 0.2 0.3 79% 94% 110% 104% 44% 49% 94.9
‘Wh-’ relatives as object 0.3 0.3 20% 13% 8% 13% 61% 62% 62.5
Table 4: Most overrepresented words in LLM texts Rates relative to human chunk 2
Llama 3 8B Llama 3 8B Instruct Llama 3 70B Llama 3 70B Instruct GPT-4o Mini GPT-4o
Word Rate Word Rate Word Rate Word Rate Word Rate Word Rate
deborah 52 unease 101 bananas 31 unease 63 camaraderie 171 camaraderie 162
rambo 22 continuation 52 paperback 30 palpable 47 tapestry 147 tapestry 155
matty 20 palpable 48 bam 26 continuation 29 palpable 145 intricate 119
goodnight 18 reminder 33 verona 25 shoutout 28 grapple 131 underscore 107
ml 15 pang 29 filth 19 intricate 27 intricate 129 unspoken 102
merlin 13 rut 29 rekall 17 pang 25 fleeting 124 amidst 100
worcester 11 waft 28 denis 14 camaraderie 24 ignite 122 palpable 95
fay 10 prioritize 27 darry 12 policymaker 24 vibrant 92 solace 95
missy 10 grapple 24 ebook 12 prioritize 24 amidst 90 fleeting 84
elisa 10 camaraderie 23 janice 12 reminder 24 cacophony 89 unravel 83
Table 5: Most underrepresented words in LLM texts Rates relative to human chunk 2
Llama 3 8B Llama 3 8B Instruct Llama 3 70B Llama 3 70B Instruct GPT-4o Mini GPT-4o
Word Rate Word Rate Word Rate Word Rate Word Rate Word Rate
bingo 0.033 anyways 0.0097 jeez 0.040 yep 0.0075 extremely 0.0034 i.e. 0.0026
scorch 0.040 analyse 0.011 donnie 0.042 fucking 0.0075 spokesman 0.0059 blah 0.0063
been 0.043 somebody 0.012 alexithymia 0.046 horrible 0.0077 i.e. 0.0095 fuck 0.0079
abt 0.043 characterise 0.013 frampton 0.048 ok 0.011 unhappy 0.010 fucking 0.0083
bananas 0.043 fuck 0.014 y’all 0.048 fuckin 0.013 bitch 0.012 asshole 0.0093
pham 0.043 obviously 0.014 analogous 0.053 amid 0.015 fucking 0.012 nasty 0.011
ie 0.045 ok 0.014 carotenoid 0.053 fortunately 0.017 kg 0.013 chum 0.013
monstrous 0.045 fuckin 0.014 mussel 0.053 i.e. 0.018 fuck 0.014 shit 0.014
pow 0.045 blah 0.015 que 0.053 obviously 0.018 visa 0.015 ok 0.014
unquote 0.045 yep 0.016 whatsapp 0.053 amongst 0.019 ter 0.016 and/or 0.016
Table 6: Word counts in the HAP-E corpus, by source
Author acad (n = 1227) blog (n = 1526) fic (n = 1395) news (n = 1322) spok (n = 1721) tvm (n = 1099) Total
Human
       Chunk 1 602,878 748,901 674,437 644,896 833,699 598,157 4,102,968
       Chunk 2 602,683 749,052 675,595 646,030 832,797 591,704 4,097,861
ChatGPT
       GPT-4o 633,884 826,590 759,461 708,375 969,210 606,059 4,503,579
       GPT-4o Mini 690,828 916,433 837,527 771,324 1,036,125 678,429 4,930,666
Llama Base
       Llama 3 70B 510,942 731,889 651,557 572,423 1,098,698 502,012 4,067,521
       Llama 3 8B 544,845 735,068 740,610 510,671 1,083,953 525,037 4,140,184
Llama Instruction-Tuned
       Llama 3 70B Instruct 580,579 735,501 641,982 627,299 934,911 453,559 3,973,831
       Llama 3 8B Instruct 5550,463 679,884 599,000 568,703 836,531 448,874 3,683,455
Total 4,717,102 6,123,318 5,580,169 5,049,721 7,625,924 4,403,831 33,500,065
Table 7: Word counts in the CAP corpus, by source
Author acad (n = 1221) blog (n = 1071) fic (n = 1372) mag (n = 1293) news (n = 1188) spok (n = 1334) tvm (n = 1006) web (n = 1130) Total
Human
       Chunk 1 594,011 521,050 669,470 629,726 579,755 651,956 497,733 549,341 4,693,042
       Chunk 2 593,875 521,493 671,400 630,126 580,208 651,013 497,900 549,736 4,695,751
ChatGPT
       GPT-4o 646,061 578,298 754,715 696,988 640,690 734,320 562,464 609,670 5,223,206
       GPT-4o Mini 710,560 631,028 830,929 759,785 697,919 793,067 608,491 664,693 5,696,472
Llama Base
       Llama 3 70B 634,881 517,613 739,502 681,171 583,279 747,996 464,058 570,922 4,939,422
       Llama 3 8B 625,065 527,628 729,348 686,137 553,703 736,093 462,948 550,569 4,871,491
Llama Instruction-Tuned
       Llama 3 70B Instruct 607,677 525,063 637,822 632,915 561,171 694,292 406,103 549,999 4,615,042
       Llama 3 8B Instruct 554,642 480,391 597,560 572,409 514,765 624,254 409,802 508,688 4,262,511
Total 4,966,772 4,302,564 5,630,746 5,289,257 4,711,490 5,632,991 3,909,499 4,553,618 38,996,937
Table 8: Random forest pairwise classification accuracy, distinguishing each LLM from human text. Training accuracy based on out-of-bag error
Trained on HAP-E Trained on CAP
LLM Training Test on CAP Training Test on HAP-E
GPT-4o
       GPT-4o 98.3% 98.4% 98.9% 97.4%
       GPT-4o Mini 98.7% 98.2% 99.2% 97.5%
Llama Instruct
       Llama 3 70B Instruct 94.5% 93.5% 94.6% 91.8%
       Llama 3 8B Instruct 95.5% 95.3% 96.1% 94.1%
Llama Base
       Llama 3 70B 93.0% 94.6% 95.8% 89.1%
       Llama 3 8B 94.0% 93.6% 95.6% 89.4%
Table 9: Lasso pairwise classification accuracy, distinguishing each LLM from human text
Trained on CAP Trained on HAP-E
LLM Training Test on HAP-E Training Test on CAP
GPT-4o
       GPT-4o 95.9% 95.3% 96.5% 95.3%
       GPT-4o Mini 96.7% 96.3% 96.9% 96.2%
Llama Instruct
       Llama 3 70B Instruct 91.1% 90.7% 92.1% 90.2%
       Llama 3 8B Instruct 92.4% 91.8% 92.9% 91.9%
Llama Base
       Llama 3 70B 77.2% 73.2% 75.6% 75.3%
       Llama 3 8B 77.6% 73.8% 76.3% 75.3%
Table 10: Random forest accuracy when trained on HAP-E and classifying M4 arXiv data.
Model Acc. (in-sample) Acc. (out-of-sample)
ChatGPT
       GPT 4o 98.22% 70.63%
       GPT 4o Mini 98.73% 60.77%
Llama Instruction-Tuned
       Llama 3 70B Instruct 95.33% 68.20%
       Llama 3 8B Instruct 96.20% 65.60%
Llama Base
       Llama 3 70B 92.43% 51.17%
       Llama 3 8B 93.56% 51.85%
Table 11: Random forest accuracy when trained on CAP and classifying M4 arXiv data.
Model Acc. (in-sample) Acc. (out-of-sample)
ChatGPT
       GPT 4o 98.76% 70.68%
       GPT 4o Mini 99.04% 65.27%
Llama Instruction-Tuned
       Llama 3 70B Instruct 95.35% 59.27%
       Llama 3 8B Instruct 96.60% 57.95%
Llama Base
       Llama 3 70B 95.53% 50.92%
       Llama 3 8B 95.39% 51.48%

### v2

@@ Results
Generalization across corpora
When each pairwise random forest was used to classify arXiv preprints from the M4 corpus, accuracy dropped significantly. Random forests trained on instruction-tuned LLMs were able to classify M4’s GPT-3.5 output with greater-than-chance accuracy, but models trained on the Llama base outputs attained only 50% accuracy, equal to random guessing (SI Appendix Tables S11–S12). These results demonstrate that instruction-tuned LLMs do have features in common that permit their classification, but that generalizability across LLMs or to different registers of text is difficult.
Discussion
This study identifies salient differences both between human and LLM-generated texts and among various models. The features that distinguish between humans and different LLMs include present participial clauses, ‘that’ clauses as sentence subject, passive voice, and nominalizations, to name a few. These findings corroborate other research that points out the ways these features produce informationally dense prose (28). Other research links these features to increased lexical diversity in generated text, as well as human judgments of linguistic mastery (13). Lastly, prior work found that ChatGPT-4 text evidences more nominalizations, and fewer human subjects and epistemic stance markers (29), findings we see reproduced in our list of distinguishing features.
A second major finding of this research is the apparently central role of instruction tuning in creating these discrepancies between human and model general texts. While we do not have access to untuned versions of GPT, comparisons between Llama’s base and tuned models emphasize the degree to which instruction tuning pushes models to produce text that reads unlike a human. This suggests that differences in style are not simply due to the selection of texts for training the base models, but due to the instruction-tuning process. Similarly, differences between GPT-4o and the instruction-tuned Llama variants may be due to differences in instruction tuning, either through different human preferences in rating responses or differences in the tasks (such as summarization) used to tune the models. (As the instruction tuning processes are not publicly documented, it is not possible to determine the cause more precisely.) While instruction tuning has previously been shown to introduce cognitive biases (30), to our knowledge, these changes in writing style are not discussed elsewhere in similar research.
@@ Discussion
As LLMs are increasingly put to work completing diverse writing tasks, these results suggest a notable misalignment between generated texts and the contexts in which we put them to use. This is another way of saying that LLMs do not vary their linguistic output in response to contextual factors in ways similar to humans. This misalignment affects experts and learners differently. For those proficient in a genre—think, a therapist collating notes on a patient or, say, a college graduate writing thank-you notes to friends and family—this misalignment is likely flagged and the output is appropriately revised. When the writer is proficient in the genre, previous experience guides a current sense of what a particular document should look like in order to be successful. For experts, then, LLMs appear a worthwhile productivity tool, suitable so long as they lend their expertise to further shaping the output.
For learners, though, LLMs appear more problematic. Of course, using LLMs to learn more about a concept, or help generate ideas, is one thing. Using output in a text is a different matter, one that may affect a student’s learning trajectories. In this case, students offload the important cognitive labor of shaping a text for a particular audience and purpose to the LLM. Never mind that LLMs do not appear to write like humans—when students offload writing work, they offload opportunities to learn how to write IMRaD articles, client reports, executive summaries, investment pitches, etc. When LLMs are used in the classroom as writing tools, instructors of all levels and disciplines need to help students see both the shortcomings of the generated text and avenues for improvement. LLMs are not bad, either technically or morally—it is only that instructors must help inculcate in students the critical perspective of the expert to know what’s working and what isn’t.
@@
[…]
[4]